How many PopTarts? are needed to convert 1,000.0 g of water at 20.0 oC to steam at 100.0 oC? One PopTart is equal to 800,000 J of energy...wow!

Grade:12th pass

1 Answers

Aziz Alam IIT Roorkee
9 years ago
Specific Heat of water= 4200 J/kg-K
Laten Heat of Vaporization=2260 J/kg

Q_n_e_t= Q_h_e_a_t_i_n_g + Q_v_a_p_o_r_i_z_i_n_g
\Rightarrow Q_n_e_t=mc\Delta T + mc_l_a_t_e_n_t
\Rightarrow Q_n_e_t=4200*80 +2260=338260 J
no of pop tarts needed are= 338260/800000=0.42285
close to half